In 6 year my age will be four time my age 9 year ago.
How old am I now?

Answer by oberobic(2304)

x = current age
.
in 6 years... x+6
.
9 years ago... x-9
.
x+6 = 4*(x-9)
.
x+6 = 4x -36
.
3x = 42
.
x = 14
.
Check this answer
.
in 6 years, 14+6 = 20
9 years ago, 14-9 = 5
4*5 = 20
Correct.
.
Answer: Current age is 14.
.
Done.